# Topology and Algebra of Bonded Knots and Braids

### v1

@@ 4.2. Rigid Standard Bonded Links: Tangle Insertions
  • Into each band, insert a properly embedded $2$–tangle chosen from a fixed family of tangles.
@@
We now state the following result, also due to Kauffman:
Theorem 4.2 (Kauffman).
@@
  • For any fixed choice of tangles, the ambient (resp. regular) isotopy class of the resulting classical link $L_{(K,B)}$ is a rigid vertex (resp. rigid vertex regular) isotopy invariant of $(K,B)$.
@@ 4.3. Rigid Tight Bonded Links: The Bonded Bracket Polynomial
In this subsection we present the bonded bracket polynomial, which extends the Kauffman bracket polynomial to rigid tight bonded links by assigning specific skein-theoretic relations at bonds. This invariant arises by choosing a particular $2$–tangle insertion at each bond, corresponding to a fixed linear combination of possible $2$–tangle states, weighted by formal coefficients. The resulting classical diagram (obtained after resolving all bonds via this prescribed $2$–tangle insertion) is then evaluated using the usual Kauffman bracket polynomial.
We emphasize that the construction described below uses one specific choice of $2$–tangle insertion at each bond. In principle, different choices of $2$–tangle insertions give rise to different invariants. Indeed, one can think of this invariant as a representative example of a much broader class of invariants defined by varying the $2$–tangle substitutions at bonds. Exploring how powerful this family of invariants is remains an open problem.
@@
The proof follows by induction, as illustrated in Figures 35 and 36, taking into account that the addition of a local curl to a bonded diagram by an R1 move multiplies the bracket polynomial of the diagram by a factor of $-A^{3}$ if the curl is right-handed and $-A^{-3}$ if the curl is left-handed.
@@
We prove relations (1) by induction on $\0n\in\mathbb{N}\backslash\{0\}$. For $n=1$ we have that:
@@ 5. Bonded Braids
In this section we develop the theory of bonded braids, an algebraic framework that extends classical braids by including bonds. Note that we restrict our attention to the standard and tight categories of bonded links, undergoing topological vertex isotopy.
We recall that a geometric braid on $n$ strands is a homeomorphic image of $n$ arcs in the interior of $[0,1]\times[0,\epsilon]\times[0,1]$, $\epsilon>0$, such that it is monotonous, that is, there are no local maxima or minima, and the boundary of the image consists in $n$ numbered points in $[0,1]\times[0,\epsilon]\times\{0\}$ and $n$ corresponding numbered points in $[0,1]\times[0,\epsilon]\times\{1\}$. We study braids through their diagrams in the plane $[0,1]\times\{0\}\times[0,1]$, which are also called braids. Formally, a braid is described by its braid word in the braid group $B_{n}$, which has generators $\sigma_{1},\sigma_{2},\ldots,\sigma_{n-1}$ and relations:
@@ 5.1. Bonded braid definition
A bonded braid on $n$ strands is a pair $(\beta,B)$ of a classical braid $\beta$ on $n$ strands, and a set of $k$ disjoint, embedded horizontal simple arcs, called bonds. The boundary points of a bond, called nodes, cannot coincide with endpoints of the braid or with other nodes, and they have local neighborhoods that are 3-valent graphs with the attaching arcs, like $\vdash$, forming together an H-neighborhood, as depicted in Figure 20. A bond joining the $i^{th}$ and $j^{th}$ strands with $i<j$ is denoted $b_{i,j}$ and it threads transversely through the strands in between. By abuse of notation, we denote by $b_{i,j}$ a bond joining the $i^{th}$ and $j^{th}$ strands with any sequence of overpasses or underpasses. A bond $b_{i,i+1}$ joining two consecutive strands $i$ and $i+1$ shall be called elementary bond, and will be denoted by $b_{i}$. If $B=\emptyset$, then $(b,\emptyset)$ is just a classical braid. Moreover, a bonded braid diagram (also referred to as bonded braid) is a regular projection of a bonded braid in the plane $[0,1]\times\{0\}\times[0,1]$ with over/under information at every crossing, and such that no crossing is horizontally aligned with a bond. So, $b_{i,j}$ can be encoded unambiguously by a sequence of $j-i+1$ o’s and u’s indicating the types of crossings formed between the bond and any strand in between. For an example see left-hand side illustration of Figure 39. A configuration of a bond which passes either over or under all its intermediate strands, such that all crossings are marked ‘over’/‘under’, is a uniform over/under.
@@ 5.2. Bonded braid isotopy
    Figure 42. Interactions between two bonds: Uniform over/under configuration.
  • Furthermore, two bonds $b_{i,j}$ and $b_{k,l}$ commute if $i<k<l<j$, the sequence of crossings of $b_{k,l}$ coincides with a subsequence of the sequence of crossings of $b_{i,j}$ for the strands $k,l$ and the markings in the bigger sequence before and after the subsequence agree, i.e. they are both ‘over’ or both ‘under’. For example, if $b_{k,l}$ forms the sequence $(o,o,u,o)$ with its intermediate strands, the bond $b_{i,j}$ must contain either the subsequence $[o,(o,o,u,o),o]$ or the subsequence $[u,(o,o,u,o),u]$ for the bonds to commute. We call this configuration the matching crossing sequences. For an illustration see Figure 43.
@@
  • Figures 48, 49 illustrate the cases where one node of the bond lies on a strand, resp. on two stands, that crosses (resp. cross) another strand. In the instance of Figure 49 there may be other strands threading through the bond, as long as they all lie under the overcrossing strand.
@@
The proof of the Lemma is adequately described in Figure 50.
@@
The additional moves depicted in Figures 42, 43 and 47 are essential to define bonded braid equivalence. These moves account for the interactions specific to bonds, ensuring that the topological properties of the bonded braid are preserved. Using these moves, we extend classical braid isotopy to include bonded braid isotopy, leading to the following formal definition.
@@ 6. The analogue of the Alexander Theorem for bonded links
In this section we focus on the theory of topological standard bonded links and we present a braiding algorithm for oriented standard bonded links. Namely, we prove the following theorem:
@@
Step 2: We next deal with up-arcs in the diagram that contain at least one node of a bond. For this we apply isotopy (TVT or RVT) as shown in Figures 57 and 58, according to whether the two joining arcs are antiparallel or parallel upward arcs. Then the bonds will only lie on down-arcs, so that the braiding algorithm will not affect them.
@@ 7. Bonded braid equivalences
Furthermore, a resolved $L$-move on a tight bonded braid is defined as above, with the addition that, if the new strands cross a bond, we pull them away using the braided vertex slide moves as described in Figure 52, see also Figure 60. Resolved $L$-moves include $L$-moves that do not cross bonds.
$L$-moves together with bonded braid isotopy generate an equivalence relation in the set of all bonded braids, the $L$-equivalence.
@@
Figure 60. A crossing formed after an $L$-move is performed.
@@
Furthermore, two oriented topological tight bonded links are topologically isotopic if and only if for any two corresponding tight bonded braids one can be obtained from the other by a finite sequence of the following moves and bonded braid isotopy:
$\begin{array}[]{lllcll}1.&{Resolved\ L-moves}&&&&\\ 2.&{Elementary\ Bond\ Commuting:}&\alpha\,b_{i}&\sim&b_{i}\,\alpha,&{for}\ \alpha,\,b_{i}\in BB_{n}\\ \end{array}$
@@
We finally check the vertex slide (VS) moves. The three representing cases of oriented VS moves with the middle arc being an up-arc are illustrated in Figures 66, 67 and 68. The moves are considered local so that all other braiding is done and we can compare the final braids. Figure 66 shows the case of parallel down-arcs at the bonding sites. After the performance of the $L$-moves we see that the move is invisible on the bonded braid level.
@@
Note that, if some arcs cross the bonds in either one of the three cases, these can be pulled away in both diagrams involved in the move, using the same vertex slide moves, so they will be braided identically. In the course of proving this Theorem it becomes apparent that it suffices to assume that all bonds are contracted to tight bonds. Likewise all $L$-moves can be assumed to be resolved $L$-moves by bonded braid isotopy. From the proof above it follows that, eventually, we only need to check the moves of Theorem 3.3.
@@ 8. The theory of enhanced bonded links & braids
We now extend the bonded knot theory by introducing enhanced bonds. In an enhanced bonded link, each bond is assigned one of two types, conventionally termed attracting vs. repelling bonds (see Figure 70). These types are meant to model different physical interactions: for example, an attracting bond might represent a bond arising from an attractive force (like a disulfide bond pulling two parts of a protein together), whereas a repelling bond may model an effective constraint that prevents two regions from approaching too closely, for example due to steric hindrance or electrostatic repulsion. Topologically, we can regard an attracting or a repelling bond as the same kind of embedded arc as before, but we treat them as ‘inverses’ of one another in an algebraic sense. We will formalize this by allowing bonds to be cancelling inverses on the level of enhanced bonded braids. In this section, we first define enhanced bonded links and the moves that relate them. Then we introduce the enhanced bonded braid group, denoted $EB_{n}$, which extends the bonded braid monoid by adding inverses for the bond generators. We then describe analogues of the Alexander and Markov theorems and $L$-equivalence in the enhanced context. We believe that the theory of enhanced bonded links can serve as a better model for biological knotted objects, as well as in other physical situations.
@@
An enhanced bonded link diagram is a bonded link diagram in which each bond behaves as in the classical bonded link setting, connecting two distinct points of $L$, but is additionally equipped with a specified interaction, giving rise to two types of bonds: attracting and repelling (see Figure 70).
@@
Note that the isotopy moves we introduced for bonded links do not alter the types assigned to the bonds in enhanced bonded links. So, enhanced bonded isotopy is defined in the same way as bonded isotopy, with the additional requirement that the bond type (attracting or repelling) assigned to each bond remains fixed throughout the isotopy. Both isotopy categories, topological and rigid vertex, continue to apply in this setting, as do the three diagrammatic forms: long, standard, and tight enhanced bonded links.
@@ 9.1. Bonded knotoids
Knotoids were introduced by Turaev in [31] as a generalization of 1-1 tangles by allowing the endpoints to be in different regions of the diagram. Namely: a knotoid diagram $K$ in an oriented surface $\Sigma$ is a generic immersion of the unit interval $[0,1]$ into $\Sigma$ whose only singularities are transversal double points endowed with over/undercrossing data called crossings. The images of 00 and $1$ under this immersion are called the endpoints of $K$ (leg and head of $K$ respectively) and are distinct from each other and from the double points (see Figure 74 by ignoring the bonds). A knotoid in $\Sigma$ is then an equivalence class of knotoid diagrams in $\Sigma$ up to the equivalence relation induced by the Reidemeister moves (recall Figures 36), that take place away from the endpoints of the knotoid.
Bonded knotoids were introduced in [11] as classical knotoids equipped with standard bonds. Here we we extend the definition to the case of having also long bonds. More precisely:
@@
A bonded knotoid diagram in $\Sigma$ is a knotoid equipped with bonds, defind as in the case of bonded knots, such that the nodes of the bonds differ from the endpoints of the knotoid. A bonded multi-knotoid diagram is defined to be a union of a bonded knotoid diagram and a finite number of knot diagrams equipped with bonds and a bonded linkoid diagram is defined to be an immersion of a disjoint union of finitely many unit intervals whose images are knotoid diagrams equipped with bonds (see Figure 74).
Note that the bonds in a bonded knotoid diagram could also be long bonds.
As in the case of bonded links, bonded knotoid diagrams in $\Sigma$ are liable to two types of equivalence relations: the topological vertex equivalence and the rigid vertex equivalence, induced by the moves of topological vertex isotopy resp. rigid vertex isotopy (recall Proposition 1.4), all taking place away from the endpoints of the bonded knotoid.
@@
An equivalence class of bonded knotoid diagrams in $\Sigma$ up to topological vertex equivalence is called a topological bonded knotoid and up to rigid vertex equivalence it is called a rigid bonded knotoid.
Figure 74. A bonded knotoid (a) and a bonded multi-knotoid (b).
Remark 9.3.
Note that in the theory of knotoids, the pulling of a strand that is adjacent to an endpoint is not allowed, since this would result into trivial knotoid diagrams. These moves are illustrated in Figure 75, and they are called forbidden moves of knotoids. Moreover, in the theory of bonded knotoids, the same is true when instead of a classical strand we consider a bond (see Figure 76). We call these moves bonded forbidden moves.
Figure 75. The forbidden moves.
@@
9.2. The bonded semi-closures of bonded knotoids
As noted in [31], the theory of knotoid diagrams suggests a new diagrammatic approach to knots, since every knotoid diagram determines a classical knot if we connect the endpoints of a knotoid diagram with an extra arc. Similarly, for bonded knotoids we define the following closure operations:

### v2

@@ 4.2. Rigid Standard Bonded Links: Tangle Insertions
  • Into each band, insert a properly embedded 222–tangle chosen from a fixed family of tangles.
@@
We now state the following result about tangle insertion, also due to Kauffman [24] but adapted to the setting of rigid bonded links. Originally the tangle insertion would take place in the disc of a rigid vertex. Here the rigid vertex is replaced by a bond. The earliest example of tangle insertion is the replacement of a singular crossing by classical crossings in the theory of Vassiliev invariants.
Theorem 4.6.
@@
[…]
@@ 4.3. Rigid Tight Bonded Links: The Bonded Bracket Polynomial
In this subsection we present the bonded bracket polynomial, which extends the Kauffman bracket polynomial to rigid tight bonded links by assigning specific skein-theoretic relations at bonds. This invariant arises by choosing a particular 222–tangle insertion at each bond, corresponding to a fixed linear combination of possible 222–tangle states, weighted by formal coefficients. The resulting classical diagram (obtained after resolving all bonds via this prescribed 222–tangle insertion) is then evaluated using the usual Kauffman bracket polynomial.
We emphasize that the construction described below uses one specific choice of 222–tangle insertion at each bond. In principle, different choices of 222–tangle insertions give rise to different invariants. Indeed, one can think of this invariant as a representative example of a much broader class of invariants defined by varying the 222–tangle substitutions at bonds. Exploring how powerful this family of invariants is remains an open problem.
@@
The proof follows by induction, as illustrated in Figures 40 and 41, taking into account that the addition of a local curl to a bonded diagram by an R1 move multiplies the bracket polynomial of the diagram by a factor of $-A^{3}$ if the curl is right-handed and $-A^{-3}$ if the curl is left-handed.
@@
We prove relations (1) by induction on $n\in\mathbb{N}\backslash\{0\}$. For $n=1$ we have that:
@@ 5. Bonded Braids
In this section we develop the theory of bonded braids, an algebraic framework that extends classical braids by including bonds. Note that we restrict our attention to the standard and tight categories of bonded links, undergoing topological vertex isotopy. Combined with the braiding result of next section, this algebraic structure could be used in encoding the topological structures of bonded links and of objects that they model.
We recall that a geometric braid on $n$ strands is a homeomorphic image of $n$ arcs in the interior of $[0,1]\times[0,\epsilon]\times[0,1]$, $\epsilon>0$, such that it is monotonous, that is, there are no local maxima or minima, and the boundary of the image consists in $n$ numbered points in $[0,1]\times[0,\epsilon]\times\{0\}$ and $n$ corresponding numbered points in $[0,1]\times[0,\epsilon]\times\{1\}$. We study braids through their diagrams in the plane $\}\times[0,1]$ 0 } × [ 0 , 1 ], which are also called braids. Formally, a braid is described by its braid word in the braid group $B_{n}$, which has generators $\sigma_{1},\sigma_{2},\ldots,\sigma_{n-1}$ and relations:
@@ 5.1. Bonded braid definition
A bonded braid on $n$ strands is a pair $(\beta,B)$ of a classical braid $\beta$ on $n$ strands, and a set of $k$ disjoint, embedded horizontal simple arcs, called bonds. The boundary points of a bond, called nodes, cannot coincide with endpoints of the braid or with other nodes, and they have local neighborhoods that are 3-valent graphs with the attaching arcs, like $\vdash$, forming together an H-neighborhood, as depicted in Figure 23. A bond joining the $i^{th}$ and $j^{th}$ strands with $i<j$ is denoted $b_{i,j}$ and it threads transversely through the strands in between. By abuse of notation, we denote by $b_{i,j}$ a bond joining the $i^{th}$ and $j^{th}$ strands with any sequence of overpasses or underpasses. A bond $b_{i,i+1}$ joining two consecutive strands $i$ and $i+1$ shall be called elementary bond, and will be denoted by $b_{i}$. If $B=\emptyset$, then $(b,\emptyset)$ is just a classical braid. Moreover, a bonded braid diagram (also referred to as bonded braid) is a regular projection of a bonded braid in the plane $\}\times[0,1]$ 0 } × [ 0 , 1 ] with over/under information at every crossing, and such that no crossing is horizontally aligned with a bond. So, $b_{i,j}$ can be encoded unambiguously by a sequence of $j-i+1$ o’s and u’s indicating the types of crossings formed between the bond and any strand in between. For an example see left-hand side illustration of Figure 44. A configuration of a bond which passes either over or under all its intermediate strands, such that all crossings are marked ‘over’/‘under’, is a uniform over/under.
@@ 5.2. Bonded braid isotopy
    Figure 47. Interactions between two bonds: uniform over configurations.
[…]
@@
  • Figures 53, 54 illustrate the cases where one node of the bond lies on a strand, resp. on two stands, that crosses (resp. cross) another strand. In the instance of Figure 54 there may be other strands threading through the bond, as long as they all lie under the overcrossing strand.
@@
The proof of the Lemma is adequately described in Figure 55. ∎
@@
The additional moves depicted in Figures 47, 48 and 52 are essential to define bonded braid equivalence. These moves account for the interactions specific to bonds, ensuring that the topological properties of the bonded braid are preserved. Using these moves, we extend classical braid isotopy to include bonded braid isotopy, leading to the following formal definition.
@@ 6. The analogue of the Alexander Theorem for bonded links
In this section we focus on the theory of topological standard bonded links and we present a braiding algorithm for oriented standard bonded links.
Namely, we prove the following theorem:
@@
Step 2: We next deal with up-arcs in the diagram that contain at least one node of a bond. For this we apply isotopy (TVT or RVT) as shown in Figures 62 and 63, according to whether the two joining arcs are antiparallel or parallel upward arcs. Then the bonds will only lie on down-arcs, so that the braiding algorithm will not affect them.
@@ 7. Bonded braid equivalences
Furthermore, a tight $L$-move on a tight bonded braid is defined as above, with the addition that, if the new strands are vertically aligned with a bond, the strands can be pulled to the side, right or left, using the braided vertex slide moves as described in Figure 57, see also Figure 65, so as to remain in the tight category. A pulled away $L$-move in the tight category shall be called tight $L$-move. Tight $L$-moves include $L$-moves that do not cross bonds.
$L$-moves (resp. tight $L$-moves) together with bonded braid isotopy generate an equivalence relation in the set of all bonded braids (resp. tight bonded braids), the $L$-equivalence (resp. tight $L$-equivalence).
@@
Figure 65. A crossing formed after an $L$-move is performed, and resolving the $L$-move by pulling strands away from the bond.
@@
Furthermore, two oriented topological tight bonded links are topologically isotopic if and only if for any two corresponding tight bonded braids one can be obtained from the other by a finite sequence of the following moves and tight bonded braid isotopy:
$\begin{array}[]{lllcll}1.&{Tight\ L-moves}&&&&\\ 2.&{Elementary\ Bond\ Commuting:}&\alpha\,b_{i}&\sim&b_{i}\,\alpha,&{for}\ \alpha,\,b_{i}\in BB_{n}\\ \end{array}$
@@
We finally check the vertex slide (VS) moves. The three representing cases of oriented VS moves with the middle arc being an up-arc are illustrated in Figures 71, 72 and 73. The moves are considered local so that all other braiding is done and we can compare the final braids. Figure 71 shows the case of parallel down-arcs at the bonding sites. After the performance of the $L$-moves we see that the move is invisible on the bonded braid level.
@@
Note that, if some arcs cross the bonds in either one of the three cases, these can be pulled away in both diagrams involved in the move, using the same vertex slide moves, so they will be braided identically. In the course of proving this Theorem it becomes apparent that it suffices to assume that all bonds are contracted to tight bonds. Likewise all $L$-moves can be assumed to be tight $L$-moves by bonded braid isotopy. From the proof above it follows that, eventually, we only need to check the moves of Theorem 3.3.
@@ 8. The theory of enhanced bonded links & braids
We now extend the bonded knot theory by introducing enhanced bonds. In an enhanced bonded link, each bond is assigned one of two types, conventionally termed attracting vs. repelling bonds (see Figure 75). These types are meant to model different physical interactions: for example, an attracting bond might represent a bond arising from an attractive force (like a disulfide bond pulling two parts of a protein together), whereas a repelling bond may model an effective constraint that prevents two regions from approaching too closely, for example due to steric hindrance or electrostatic repulsion. Topologically, we can regard an attracting or a repelling bond as the same kind of embedded arc as before, but we treat them as ‘inverses’ of one another in an algebraic sense. We will formalize this by allowing bonds to be cancelling inverses on the level of enhanced bonded braids.
In this section, we first define enhanced bonded links and the moves that relate them. Then we introduce the enhanced bonded braid group, denoted $EB_{n}$, which extends the bonded braid monoid by adding inverses for the bond generators. We then describe analogues of the Alexander and Markov theorems and $L$-equivalence in the enhanced context. We believe that the theory of enhanced bonded links can serve as a better model for biological knotted objects, as well as in other physical situations.
@@
An (oriented) enhanced bonded knot/link is a pair $(L,B)$, where $L$ is an (oriented) link in $S^{3}$, and $B=B_{+}\cup B_{-}$ is a finite set of bonds as defined in Definition 1.1 and such that each element of the subset $B_{+}$ is equipped with two arrows pointing toward one another, and each element of the subset $B_{-}$ is equipped with two arrows pointing away from one another. Call elements of $B_{+}$ attracting bonds and elements of $B_{-}$ repelling bonds. For an example view Figure 75 where the enhanced bonds are standard. Moreover, two adjacent bonds are called parallel if together they form two opposite sides of a standardly embedded rectangular band, so that the other two sides are link arcs, and such the band can be contracted by isotopy to a regularly projected planar tight rectangle (in the sense that its interior does not intersect other arcs or bonded arcs). Further, if we have two adjacent parallel bonds of opposite type, then they cancel, meaning that the pair of bonds can be removed. An enhanced bonded link diagram is a diagram of an (oriented) enhanced bonded knot/link.
@@
Regarding isotopy, the moves we introduced for bonded links do not alter the types assigned to the bonds in enhanced bonded links. So, enhanced bonded isotopy is defined in the same way as bonded isotopy, with the additional requirements that the bond type (attracting or repelling) assigned to each bond remains fixed throughout the isotopy, and that parallel bonds of opposite type cancel, view Figure 76. Both isotopy categories, topological and rigid vertex, continue to apply in this setting, as do the three diagrammatic forms: long, standard, and tight enhanced bonded links.
Figure 76. Cancellation of parallel bonds of opposite type.
Remark 8.2.
Note that it is possible to generalize the definition of enhanced bonded link diagrams to include bonds that are without the above form of enhancement (so that they do not have inverses) and also one can include many types of bonds, with or without inverses, if these are needed. We retain the above definition in our discussion for the sake of simplicity.
@@ 9.1. Bonded knotoids
Knotoids were introduced by Turaev in [41] as a generalization of 1-1 tangles by allowing the endpoints to be in different regions of the diagram. Namely: a knotoid diagram $K$ in an oriented surface $\Sigma$ is a generic immersion of the unit interval [0,1][0,1][ 0 , 1 ] into $\Sigma$ whose only singularities are transversal double points endowed with over/undercrossing data called crossings. The images of 0 and 111 under this immersion are called the endpoints of $K$ (leg and head of $K$ respectively) and are distinct from each other and from the double points (see Figure 81 by ignoring the bonds).
A knotoid in $\Sigma$ is then an equivalence class of knotoid diagrams in $\Sigma$ up to the equivalence relation induced by the Reidemeister moves (recall Figures 58), that take place away from the endpoints of the knotoid. A multi-knotoid diagram in $\Sigma$ is a generic immersion of the unit interval [0,1][0,1][ 0 , 1 ] and a number of copies of $S^{1}$ into $\Sigma$. Note that in the theory of multi-knotoids, the pulling of a strand that is adjacent to an endpoint is not allowed, since this would result into trivial knotoid diagrams. These moves are illustrated in Figure 80, and they are called forbidden moves of multi-knotoids. A knotoid has a natural orientation from leg to head. A multi-knotoid diagram is oriented if orientations are also assigned to its closed components.
Figure 80. The forbidden moves.
Bonded knotoids were introduced in [15] as classical knotoids equipped with standard bonds. Here we we extend the definition to the case of multi-knotoids having also long bonds. More precisely:
@@
[…]
Figure 81. A bonded knotoid (a) and a bonded multi-knotoid (b).
As in the case of bonded links, bonded multi-knotoid diagrams in $\Sigma$ are liable to two types of equivalence relations: the topological vertex equivalence and the rigid vertex equivalence, induced by the moves of topological vertex isotopy resp. rigid vertex isotopy (recall Proposition 1.5), all taking place away from the endpoints of the bonded multi-knotoid.
@@
An equivalence class of bonded multi-knotoid diagrams in $\Sigma$ up to topological vertex equivalence is called a topological bonded multi-knotoid and up to rigid vertex equivalence it is called a rigid bonded multi-knotoid.
In the theory of bonded knotoids w have the forbidden moves for the multi-knotoid arcs as well as for the bonds (see Figure 82). We call these moves bonded forbidden moves.
@@
9.2. Types of closures for bonded multi-knotoids
As noted in [41], the theory of knotoids suggests a new diagrammatic approach to the theory of knots, since a knotoid diagram gives rise to a classical knot if we connect its endpoints with an extra simple arc. Similarly, for bonded multi-knotoids we define the following closure operations: